OpenStudy (anonymous):

Natasha drove from Bedingfield to Portsmouth at an average speed of 100km/h. On her way back, she drove 75km/h. Her trip involved a total of 3.5hrs of driving time. What is the distance between Bedingfield and Portsmouth? I just need the equation to set this up.

OpenStudy (anonymous):

speed x time =distance

OpenStudy (anonymous):

yeah but 3.5 hours is a combination of the time on the way over and back, and there are two different speeds

OpenStudy (anonymous):

time going is \[\frac{D}{100}\] time returning is \[\frac{D}{75}\] the total time is 3.5 so put \[\frac{D}{100}+\frac{D}{75}=3.5\] solve for D

OpenStudy (anonymous):

that is the equation set up. you can solve this yes?

OpenStudy (anonymous):

yeah just let me work it out really quick. the answer should be 150

OpenStudy (anonymous):

ok. work it out and i bet that is what you get

OpenStudy (anonymous):

its obviously 150 i just can't ever think of how to set it up

OpenStudy (anonymous):

yep! thanks

OpenStudy (anonymous):

sure is i just checked! yw